OpenStudy (anonymous):

The area of a rectangle is 12 square inches. The length is 5 more than twice the width. Find the length and the width

OpenStudy (anonymous):

Let the width equal x, therefore the length is 2x+5. The area is length*width, so we know: x(2x+5)=12 => 2x^2+5x=12 => 2x^2+5x-12=0 => (2x-3)(x+4)=0 So we know x=-4, or x=1.5. Lengths can't be negative, so x=1.5.